\begin{thebibliography}{998}
%\bibitem{AN52}Ankeny, N. C. \textit{\color{blue}Representations of primes by quadratic forms.} Amer. J. Math. 74 (1952), 913-919. 
\bibitem{AP98} Apostol, Tom M. \textit{\color{blue}Introduction to analytic number theory.} Undergraduate Texts in Mathematics. Springer-Verlag, New York-Heidelberg, 1976.

\bibitem{BD96} Bateman, Paul T.; Diamond, Harold G. \textit{\color{blue}A hundred years of prime numbers}. Amer. Math. Monthly 103 (1996), no. 9, 729-741.
\bibitem{BF88} Bohman, Jan; Frieberg; Carl-Erik. \textit{\color{blue}The Stieltjes function definition and properties}. Math. Comp. 51 (1988), no. 183, 281-289.
%\bibitem{BJ14} Borwein, Jonathan M.\textit{\color{blue} The life of $\pi$: from Archimedes to ENIAC and beyond. From Alexandria, through Baghdad}, %531%-561, Springer, Heidelberg, 2014.
%\bibitem{BB04} Berggren, Lennart; Borwein, Jonathan; Borwein, Peter.  \textit{\color{blue}Pi: a source book}. Third edition. Springer-Verlag, New York, %2004.
%\bibitem{BB87} Borwein, J. M. and Borwein, P. B.  \textit{\color{blue}AGM: A Study in Analytic Number Theory and Computational Complexity}. New York: %Wiley, pp. 362-386, 1987.
%\bibitem{BC13} Jason Bell, Yann Bugeaud, Michael Coons. Diophantine approximation of Mahler numbers, arXiv:1307.4123.
%\bibitem{BF00} Beukers, Frits. \textit{\color{blue}A rational approach to $\pi$.}  Nieuw Arch. Wiskd. (5) 1 (2000), no. 4, 372-379.

%\bibitem{BR19} Roger Baker. \textit{\color{blue}A theorem of Bombieri-Vinogradov type with few exceptional moduli,} arXiv:1905.12488.


%\bibitem{BT19}Oleksiy Klurman, Alexander P. Mangerel, Joni Teravainen. \textit{\color{blue}Multiplicative functions in short arithmetic progressions,} arXiv:1909.12280.


%\bibitem{BY08} Bugeaud, Yann.  \textit{\color{blue}Diophantine approximation and Cantor sets}. Math. Ann. 341 (2008), no. 3, 677-684.
%\bibitem{CG82} Chudnovsky, G. V. \textit{\color{blue}Hermite-Pade approximations of the measure of irrationality of $\pi$}. Lecture Notes in Math., %925, Springer, Berlin-New York, 1982.
%\bibitem{CJ90} Conway, John B. \textit{\color{blue}A course in functional analysis.} Second edition. Graduate Texts in Mathematics, 96. Springer-%Verlag, New York, 1990.
%\bibitem{CP19} S. F. Chen, Erin P. J. Pearse. \textit{\color{blue}The irrationality measure of $\pi$ as seen through the eyes of $cos(n)$,} %arXiv:1807.02955.
%\bibitem{DLMF} NIST. \textit{\color{blue}Digital Library Mathematical Functions.}https://dlmf.nist.gov.
\bibitem{DS70} Diamond, Harold G.; Steinig, John. \textit{\color{blue}An elementary proof of the prime number theorem with a remainder term}. Invent. Math. 11 (1970), 199-258.

 
% \bibitem{KT89} Korner, T. W. \textit{\color{blue}Fourier analysis.} Second edition. Cambridge University Press, Cambridge, 1989.
%\bibitem{LS95}  Lang, Serge. \textit{\color{blue}Introduction to Diophantine approximations.} Second edition. Springer-Verlag, New York, 1995.
%\bibitem{MK53} Mahler, K. \textit{\color{blue}On the approximation of $\pi$}. Nederl. Akad. Wetensch. Proc. Ser. A. 56=Indagationes Math. 15, (1953). %30-42. 
%^\bibitem{MM74} Mignotte, M. \textit{\color{blue}Approximations rationnelles de $\pi$ et quelques autres nombres}. Journees Arithmetiques (Grenoble, %1973), pp. 121-132. Bull. Soc. Math. France, Mem. 37, 1974.
%\bibitem{NZ91} Niven, Ivan; Zuckerman, Herbert S.; Montgomery, Hugh L.  \textit{\color{blue}An introduction to the theory of numbers.} Fifth edition. %John Wiley $\&$ Sons, Inc., New York, 1991. 
%\bibitem{OC63} Olds C. E. \textit{\color{blue}Continued Fraction.} Random House, New York 1963. 
%\bibitem{OEIS} On-Line Encyclopedia of Integer Sequences, oeis.org.
%\bibitem{PC07} Clifford A. Pickover. \textit{\color{blue}The Mathematics of Oz: Mental Gymnastics from Beyond the Edge}. Cambridge University Press, %2007.
%\bibitem{RH94} Rose, H. E.  \textit{\color{blue}A course in number theory.} Second edition. Oxford Science Publications. The Clarendon Press, Oxford %University Press, New York, 1994.
%\bibitem{SJ05} Steuding, Jorn.  \textit{\color{blue}Diophantine analysis.} Discrete Mathematics and its Applications (Boca Raton). Chapman $\&$ %Hall/CRC, Boca Raton, FL, 2005.
%\bibitem{SV08}  Salikhov, V. Kh. \textit{\color{blue}On the irrationality measure of $\pi$}. Russian Math. Surveys 63 (2008), no. 3, 570-572. 
%\bibitem{TG17} Thomas, G. B. \textit{\color{blue}Thomas' Calculus: Early Transcendentals} 12th Edition, Edited by Joel R. Hass, Christopher E. Heil, et 
%al. 2017.
%\bibitem{WM00} Waldschmidt, Michel. \textit{\color{blue}Diophantine approximation on linear algebraic groups. Transcendence properties of the %exponential function in several variables.} Grundlehren der Mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], 326. %Springer-Verlag, Berlin, 2000.
%\bibitem{HM93} Hata, M. "Rational Approximations to  and Some Other Numbers." Acta Arith. 63, 335-349, 1993. 
%M. Hata. J. Reine Angew. Math., 407 (1990), 99-125.
%\bibitem{AP76} Apostol, Tom M. \textit{\color{blue}Introduction to analytic number theory.} Undergraduate Texts in Mathematics. Springer-Verlag, New %York-Heidelberg, 1976.
%\bibitem{AO83} Adleman, Leonard M.; Odlyzko, Andrew M. \textit{\color{blue}Irreducibility testing and factorization of polynomials}. Math. Comp. 41  %(1983),  no. 164, 699-709. 
%\bibitem{BA83} Balog, A. \textit{\color{blue}On the fractional part of $p^{\theta}$}. Arch. Math. (Basel)  40  (1983),  no. 5, 434-440.
%\bibitem{BA83} Balog, A. \textit{\color{blue}On the fractional part of $p^{\theta}$}. Arch. Math. (Basel) 40 (1983), no. 5, 434-440.
%\bibitem{BI89} Bombieri, E.; Friedlander, J. B.; Iwaniec, H. \textit{\color{blue}Primes in arithmetic progressions to large moduli. III.} J. Amer. Math. Soc. 2 (1989), no. 2, 215-224.
%\bibitem{BH62} P.T. Bateman, R.A. Horn. \textit{\color{blue}A heuristic asymptotic formula concerning the distribution of prime numbers.} Math. Comput., 16 (1962), pp. 363-367.
%\bibitem{BO15} Bordelles, Olivier. \textit{\color{blue}Some explicit estimates for the Mobius function.} J. Integer Seq. 18 (2015), no. 11, Article 15.11.1, 13 pp.
%\bibitem{BZ07} Stephan Baier, Liangyi Zhao. \textit{\color{blue}On Primes Represented by Quadratic Polynomials}, arXiv:math/0703284.
%\bibitem{BZ07B} Baier, Stephan; Zhao, Liangyi. \textit{\color{blue}Primes in quadratic progressions on average}. Math. Ann. 338 (2007), no. 4, 963-982. 
%\bibitem{BS09} Banks, William D.; Shparlinski, Igor E. \textit{\color{blue}Prime numbers with Beatty sequences}. Colloq. Math.  115  (2009),  no. 2, %147-157. 
%\bibitem{BR95} Baker, R. C.; Harman, G.; Rivat, J. \textit{\color{blue}Primes of the form $n^c$.} J. Number Theory  50  (1995),  no. 2, 261-277. 
%\bibitem{CC00} Chris K. Caldwell. \textit{\color{blue}An Amazing Prime Heuristic, Preprint, 2000.
%\bibitem{CA04}Cojocaru, Alina Carmen. \textit{\color{blue}Questions about the reductions modulo primes of an elliptic curve}. Number theory, 61-79, CRM Proc. Lecture Notes, 36, Amer. Math. Soc., Providence, RI, 2004.
%\bibitem{CJ11} Cilleruelo, Javier. \textit{\color{blue} The least common multiple of a quadratic sequence.} Compos. Math. 147 (2011), no. 4, 1129-1150.
%\bibitem{CM15} Giovanni Coppola, M. Ram Murty, Biswajyoti Saha. \textit{\color{blue}On the error term in a Parseval type formula in the theory of %Ramanujan expansions II, arXiv:1507.07862.
%\bibitem{CN09} Childress, Nancy Class field theory. Universitext. Springer, New York, 2009. 
%\bibitem{CO07} Cohen, Henri. \textit{\color{blue}Number theory. Vol. II. Analytic and modern tools. Graduate Texts in Mathematics, 240. Springer, New %York, 2007.
%\bibitem{DE76} Delange, Hubert. \textit{\color{blue}On Ramanujan expansions of certain arithmetical functions. Acta Arith. 31 (1976), no. 3, 259-270.
%\bibitem{DI82} J.M. Deshouillers and H. Iwaniec. \textit{\color{blue}On the greatest prime factor of $n^2 + 1$}, Ann. Inst. Fourier (Grenoble) 32 %(1982). 
%\bibitem{DH00} Davenport, Harold. \textit{\color{blue}Multiplicative number theory. Third edition. Revised and with a preface by Hugh L. Montgomery. %Graduate Texts in Mathematics, 74. Springer-Verlag, New York, 2000.
\bibitem{DS70} Diamond, Harold G.; Steinig, John. \textit{\color{blue}An elementary proof of the prime number theorem with a remainder term}. Invent. Math. 11 1970 199-258.

%\bibitem{DL12} De Koninck, Jean-Marie; Luca, Florian. \textit{\color{blue}Analytic number theory. Exploring the anatomy of integers.} Graduate Studies in Mathematics, 134. American Mathematical Society, Providence, RI, 2012.

\bibitem{DLMF} NIST \textit{\color{blue}Digital Library of Mathematical Functions}. http://dlmf.nist.gov/, 2019. F. W. J. Olver, ..., and M. A. McClain, eds.
\bibitem{EL85}  Ellison, William; Ellison, Fern. \textit{\color{blue}Prime numbers.} A Wiley-Interscience Publication. John Wiley and Sons, Inc., New %York; Hermann, Paris, 1985.
%\bibitem{EL00} Leonhard Euler.  \textit{\color{blue}De Numeris Primis Valde Magnis}, Novi Commentarii academiae scientiarum Petropolitanae 9, 1764, pp. 99-153. http://eulerarchive.maa.org/docs/originals/E283.pdf.
%\bibitem{FG18}Soren Laing Aletheia-Zomlefer, Lenny Fukshansky, Stephan Ramon Garcia. \textit{\color{blue}The Bateman-Horn Conjecture: Heuristics, History, and Applications,} arXiv:1807.08899.
\bibitem{EP49} Erdos, P. \textit{\color{blue}On a new method in elementary number theory which leads to an elementary proof of the prime number theorem}. Proc. Nat. Acad. Sci. U.S.A. 35 (1949), 374-384.
%\bibitem{EP60} Erdos, P.  \textit{\color{blue}On a problem of G. Golomb}. J. Austral. Math. Soc. 2 1961/1962 1-8. 

\bibitem{FI10} Friedlander, John; Iwaniec, Henryk. \textit{\color{blue}Opera de cribro}. American Mathematical Society Colloquium Publications, 57. American Mathematical Society, Providence, RI, 2010.
%\bibitem{FG92} Friedlander, John; Granville, Andrew. \textit{\color{blue}Limitations to the equi-distribution of primes. III}. Compositio Math. 81 (1992), no. 1, 19-32.
\bibitem{FG91}  Friedlander, John; Granville, Andrew. \textit{\color{blue}Limitations to the equi-distribution of primes. IV}. Proc. Roy. Soc. London Ser. A  435  (1991),  no. 1893, 197-204.
%\bibitem{FK02}  Ford, Kevin. \textit{\color{blue}Vinogradov's integral and bounds for the Riemann zeta function}. Proc. London Math. Soc. (3) 85 (2002), no. 3, 565-633. 

%\bibitem{HA05}  Hildebrand, A. J. \textit{\color{blue}Arithmetic Functions II: Asymptotic Estimates}, www.math.uiuc.edu/~hildebr/ant/main2.pdf.

%\bibitem{HW08} Hardy, G. H.; Wright, E. M. \textit{\color{blue}An introduction to the theory of numbers}. Sixth edition. Oxford University Press, Oxford, 2008.




\bibitem{FJ15} Friedlander, John B. \textit{\color{blue}Counting primes in arithmetic progressions}. Analytic number theory, 101-114, Springer, Cham, 2015.

%\bibitem{GK06} Greene, Robert E.; Krantz, Steven G.  \textit{\color{blue}Function theory of one complex variable}. Graduate Studies in Mathematics, 40. %American Mathematical Society, Providence, RI, 2006. 

%\bibitem{HM93} Hata, Masayoshi.  \textit{\color{blue}A lower bound for rational approximations to $\pi$}. J. Number Theory 43 (1993), no. 1, 51-67.
%\bibitem{HW08} G. H. Hardy and E. M. Wright.  \textit{\color{blue}An Introduction to the Theory of Numbers.} 5th ed., Oxford University Press, Oxford, %2008.




%\bibitem{FK12} Christopher F. Fowler, Stephan Ramon Garcia, Gizem Karaali. \textit{\color{blue}Ramanujan sums as supercharacters, arXiv:1201.1060.
%\bibitem{BS18} Olivier Bordelles, Randell Heyman, Igor E. Shparlinski. \textit{\color{blue}On a sum involving the Euler function}, arXiv:1808.00188.
%\bibitem{GL10} Luca Goldoni. \textit{\color{blue}Prime Numbers And Polynomials}, Phd Thesis, Universita` Degli Studi Di Trento, 2010.
%\bibitem{FZ18}Soren Laing Aletheia-Zomlefer, Lenny Fukshansky, Stephan Ramon Garcia.\textit{\color{blue} The Bateman-Horn Conjecture: Heuristics, %History, and Applications}. arXiv:1807.08899
\bibitem{GR09} Granville, Andrew. \textit{\color{blue}Different approaches to the distribution of primes.} Milan J. Math. 78 (2010), no. 1, 65-84.
\bibitem{GR92} Granville, A. \textit{\color{blue}On elementary proofs of the prime number theorem for arithmetic progressions, without characters}. Proc. Amalfi Conference on Analytic Number Theory, 157-194, Univ. Salerno, Salerno, 1992.
%\bibitem{GD09} D. A. Goldston. \textit{\color{blue}Are There Infinitely Many Twin Primes?, Preprint, 2009.
%\bibitem{GG05} Goldston, D. A.; Graham, S. W.; Pintz, J.; Yildirim, C. Y. \textit{\color{blue}Small gaps between almost primes, the parity problem, and %some conjectures of Erdos on consecutive integers, arXiv:math.NT/0506067.
%\bibitem{GL10} Luca Goldoni, \textit{\color{blue}Prime Numbers And Polynomials, Phd Thesis, Universita` Degli Studi Di Trento, 2010.
%\bibitem{GM00} Granville, Andrew; Mollin, Richard A. \textit{\color{blue}Rabinowitsch revisited}. Acta Arith. 96 (2000), no. 2, 139-153. 
%\bibitem{JW03} Jacobson, Michael J., Jr.; Williams, Hugh C. \textit{\color{blue}New quadratic polynomials with high densities of prime values}. Math. Comp. 72 (2003), no. 241, 499-519.
%\bibitem{GP99} Gadiyar, H. Gopalkrishna; Padma, R. \textit{\color{blue}Ramanujan-Fourier series, the Wiener-Khintchine formula and the distribution of %prime pairs. Phys. A  269  (1999),  no. 2-4, 503-510.
%\bibitem{GP06} H. Gopalkrishna Gadiyar, R. Padma. \textit{\color{blue}Linking the Circle and the Sieve: Ramanujan - Fourier Series, arXiv:math/0601574.

\bibitem{HG07} Harman, Glyn. \textit{\color{blue}Prime-detecting sieves}. London Mathematical Society Monographs Series, 33. Princeton University Press, Princeton, NJ, 2007.
%\bibitem{HG83} Harman, Glyn. \textit{\color{blue}On the distribution of $\{\sqrt{p}\}$ modulo one}. Mathematika  30  (1983),  no. 1, 104-116. 
%\bibitem{HR95} Harman, Glyn; Rivat, Joel. \textit{\color{blue}Primes of the form $[p^c ]$ and related questions.} Glasgow Math. J.  37  (1995),  no. 2, %131-141.
%\bibitem{HG16} Harman, Glyn. \textit{\color{blue}Primes in Beatty sequences in short intervals}. Mathematika  62  (2016),  no. 2, 572-586. 
%\bibitem{HL23} Hardy, G. H. Littlewood J.E. \textit{\color{blue}Some problems of Partitio numerorum III: On the expression of a number as a sum of primes}. Acta Math. 44 (1923), No. 1, 1-70.
%\bibitem{HR59} Hardy, G. H. Ramanujan: \textit{\color{blue}Twelve lectures on subjects suggested by his life and work.  Chelsea Publishing Company, New %York 1959.
%\bibitem{HG16}Hurst, Greg. \textit{\color{blue}Computations of the Mertens function and improved bounds on the Mertens conjecture}. Math. Comp. 87 %(2018), no. 310, 1013-1028 or arxiv.1610.08551.
\bibitem{HM72}Huxley, M. N. \textit{\color{blue}The distribution of prime numbers. Large sieves and zero-density theorems}. Oxford Mathematical Monographs. Clarendon Press, Oxford, 1972.
\bibitem{IK04} Iwaniec, Henryk; Kowalski, Emmanuel. \textit{\color{blue}Analytic number theory.} AMS Colloquium Publications, 53. American Mathematical Society, Providence, RI, 2004.
%\bibitem{KA97} Khinchin, A. Ya. \textit{\color{blue}Continued fractions.} Reprint of the 1964 translation. Dover Publications, Inc., Mineola, NY, 1997. 
%\bibitem{HW08} Hardy, G. H.; Wright, E. M. An introduction to the theory of numbers. Sixth edition. Revised by D. R. Heath-Brown and J. H. Silverman. %With a foreword by Andrew Wiles. Oxford University Press, Oxford, 2008.
%\bibitem{IK04} Iwaniec, Henryk; Kowalski, Emmanuel. \textit{\color{blue}Analytic number theory}. AMS Colloquium Publications, 53. American Mathematical %Society, Providence, RI, 2004. 
%\bibitem{IH78} Iwaniec, Henryk. \textit{\color{blue}Almost-primes represented by quadratic polynomials.} Invent. Math. 47 (1978), no. 2, 171-188. 
%\bibitem{JW03} Jacobson, Michael J., Jr.; Williams, Hugh C. \textit{\color{blue}New quadratic polynomials with high densities of prime values. Math. %Comp. 72 (2003), no. 241, 499-519.
%\bibitem{KA99} Kumchev, A. \textit{\color{blue}On the distribution of prime numbers of the form $[n^c ]$} . Glasg. Math. J.  41  (1999),  no. 1, 85-%102.
%\bibitem{KE05} Emmanuel Kowalski. \textit{\color{blue}Analytic problems for elliptic curves}. arXiv:math/0510197.
 

%\bibitem{KJ92} Keiper, J. B. \textit{\color{blue}Power series expansions of Riemann's $\xi$ function. Math. Comp. 58 (1992), no. 198, 765-773.
%\bibitem{KL12} Kunik, Matthias; Lucht, Lutz G. \textit{\color{blue}Power series with the von Mangoldt function. Funct. Approx. Comment. Math. 47 %(2012),  part 1, 15-33. 
%\bibitem{KR09} Jacob Korevaar. \textit{\color{blue}Prime pairs and the zeta function, Journal of Approximation Theory 158 (2009) 69-96.
%\bibitem{KT06} Kotnik, Tadej; te Riele, Herman. \textit{\color{blue}The Mertens conjecture revisited}. Algorithmic number theory, 156-167, Lecture %Notes in Comput. Sci., 4076, Springer, Berlin, 2006.
%\bibitem{LL95} Lucht, Lutz. \textit{\color{blue}Ramanujan expansions revisited. Arch. Math. (Basel) 64 (1995), no. 2, 121-128.
%\bibitem{LL10} Lucht, Lutz G. \textit{\color{blue}A survey of Ramanujan expansions. Int. J. Number Theory 6 (2010), no. 8, 1785-1799. 
%\bibitem{LR12} Lemke Oliver, Robert J. \textit{\color{blue}Almost-primes represented by quadratic polynomials.} Acta Arith. 151 (2012), no. 3, 241-261.
%\bibitem{MA09} Matomaki, Kaisa. \textit{\color{blue}A note on primes of the form $p=aq^2+1$}. Acta Arith. 137 (2009), no. 2, 133-137.
%\bibitem{MK86} McCurley, Kevin S. \textit{\color{blue}The smallest prime value of $x^n +a$} . Canad. J. Math.  38  (1986),  no. 4, 925-936.
%\bibitem{MM13} Mariusz Mirek. \textit{\color{blue}Roth Theorem in the Piatetski-Shapiro primes}, arXiv:1305.0043. 
%\bibitem{MR19} James Maynard, Zeev Rudnick. \textit{\color{blue}A lower bound on the LCM of polynomial sequences}, arXiv:1910.13218.
%\bibitem{MK84} McCurley, Kevin S. \textit{\color{blue}Prime values of polynomials and irreducibility testing}. Bull. Amer. Math. Soc. (N.S.) 11 (1984), no. 1, 155-158. 
%\bibitem{MS07} Maier, H.; Stewart, C. L. \textit{\color{blue}On intervals with few prime numbers. J. Reine Angew. Math. 608 (2007), 183-199.
\bibitem{MH88}Maier, Helmut. \textit{\color{blue}Primes in short intervals}. Michigan Math. J. 32 (1985), no. 2, 221-225. 
%\bibitem{MT06} Miller, Steven J.; Takloo-Bighash, Ramin. \textit{\color{blue}An invitation to modern number theory. With a foreword by Peter Sarnak. %Princeton University Press, Princeton, NJ, 2006. 
\bibitem{MV07} Montgomery, Hugh L.; Vaughan, Robert C. \textit{\color{blue}Multiplicative number theory. I. Classical theory.} Cambridge University %Press, Cambridge, 2007.
%\bibitem{MT06} Miller, Steven J.; Takloo-Bighash, Ramin. \textit{\color{blue}An invitation to modern number theory}. With a foreword by Peter Sarnak. %Princeton University Press, Princeton, NJ, 2006.
\bibitem{NW00} Narkiewicz, W. \textit{\color{blue}The development of prime number theory. From Euclid to Hardy and Littlewood}. Springer Monographs in Mathematics. Springer-Verlag, Berlin, 2000. 
%\bibitem{PP} Prime Pages,  www.prime.utm.edu.

%\bibitem{PS12} Paul Pollack, Carlo Sanna, Uncertainty principles connected with the Mobius inversion formula, arXiv:1211.0189. 

%\bibitem{PA88}  A. G. Postnikov. \textit{\color{blue}Introduction to analytic number theory}, Translations of Mathematical Monographs, vol. 68, American Mathematical Society, Providence, RI, 1988.



%\bibitem{SS94}  Schwarz, Wolfgang; Spilker, Jurgen. Arithmetical functions. An introduction to elementary and analytic properties of arithmetic. London Mathematical Society Lecture Note Series, 184. Cambridge University Press, Cambridge, 1994.
%\bibitem{PJ09} Pintz, Janos. \textit{\color{blue}Landau's problems on primes}. J. Theory. Nombres Bordeaux 21 (2009), no. 2, 357-404.
%\bibitem{PK57} Prachar, Karl. \textit{\color{blue} Handbuch der Lehre von der Verteilung der Primazahlen [1909]} Berlin, Springer, 1957. 
%\bibitem{RM13} Ram Murty, M. \textit{\color{blue}Ramanujan series for arithmetical functions. Hardy-Ramanujan J.  36  (2013), 21-33.
%\bibitem{RD96} Redmond, Don.  \textit{\color{blue}Number theory. An introduction}. Monographs and Textbooks in Pure and Applied Mathematics, 201. Marcel Dekker, Inc., New York, 1996. 
%\bibitem{RI15} Igor Rivin. \textit{\color{blue}Some experiments on Bateman-Horn,} arXiv:1508.07821.
%\bibitem{RP96} Ribenboim, Paulo. \textit{\color{blue}The new book of prime number records}, Berlin, New York: Springer-Verlag, 1996.
%\bibitem{RS01} Rivat, Joel; Sargos, Patrick. \textit{\color{blue}Nombres premiers de la forme $[n^ c]$.} Canad. J. Math.  53  (2001),  no. 2, 414-433.
%\bibitem{RI15} Igor Rivin. \textit{\color{blue}Some experiments on Bateman-Horn,  arXiv:1508.07821.
\bibitem{SA49} Selberg, Atle. \textit{\color{blue}An elementary proof of the prime-number theorem}. Ann. of Math. (2)  50,  (1949). 305-313.
\bibitem{SA49B} Selberg, Atle. \textit{\color{blue}An elementary proof of Dirichlet's theorem about primes in an arithmetic progression}. Ann. of Math. (2) 50 (1949), 297-304. 
%\bibitem{SH83} Shapiro, Harold N. \textit{\color{blue}Introduction to the theory of numbers}. Pure and Applied Mathematics. A Wiley-Interscience Publication. New York, 1983.
%\bibitem{SJ08} Jorn Steuding, Marc Technau. \textit{\color{blue}The Least Prime Number in a Beatty Sequence,} arXiv:1512.08382.
%\bibitem{SD60} Shanks, Daniel. \textit{\color{blue}On numbers of the form $n^4 +1$ }. Math. Comput.  15,  1961, 186-189.
%\bibitem{SD61} Shanks, Daniel. \textit{\color{blue}A sieve method for factoring numbers of the form $n^2 +1$ }. Math. Tables Aids Comput.  13,  1959, 78-86.
%\bibitem{SK07} K. Soundararajan. \textit{\color{blue}Partial sums of the Mobius function}, arXiv:0705.0723.
 \bibitem{SP73} Serre, J.-P. \textit{\color{blue}A course in arithmetic}. Graduate Texts in Mathematics, No. 7. Springer-Verlag, New York-Heidelberg, 1973.
%\bibitem{SP80} Spilker, Jurgen. \textit{\color{blue}Ramanujan expansions of bounded arithmetic functions. Arch. Math. (Basel) 35 (1980), no. 5, 451-%453.
%\bibitem{SS94} Schwarz, Wolfgang; Spilker, Jurgen. \textit{\color{blue}Arithmetical functions. An introduction to elementary and analytic properties of %arithmetic functionss. London Mathematical Society Lecture Note Series, 184. Cambridge University Press, Cambridge, 1994.
%\bibitem{TG15} Tenenbaum, Gerald. \textit{\color{blue}Introduction to analytic and probabilistic number theory. Translated from the Third French %edition. American Mathematical Society, Rhode Island, 2015.
%\bibitem{TA00} Toth, Arpad. \textit{\color{blue}Roots of quadratic congruences}. Internat. Math. Res. Notices  2000,  no. 14, 719-739. 
%\bibitem{VR97} Vaughan, R. C. \textit{\color{blue}The Hardy-Littlewood method. Second edition. Cambridge Tracts in Mathematics, 125. Math. Soc. (N.S.) %44 (2007), no. 1, 1-18. Cambridge University Press, Cambridge, 1997. 
%\bibitem{VR73} Vaughan, R. C. \textit{\color{blue}Some applications of Montgomery's sieve}. J. Number Theory 5 (1973), 64-79.
%\bibitem{WN66} Wiener, Norbert. \textit{\color{blue}Generalized harmonic analysis and Tauberian theorems. M.I.T. Press, Cambridge, Mass., London 1966.
%\end{thebibliography}








%\begin{thebibliography}{99}


%\bibitem{KJ92}  Keiper, J. B. Power series expansions of Riemann's $\xi$ function. Math. Comp. 58 (1992), no. 198, 765-773.

%\bibitem{LD88}  Lehmer, D. H. The sum of like powers of the zeros of the Riemann zeta function. Math. Comp. 50 (1988), no. 181, 265-273.




\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[top=3cm, bottom=3cm, left=3cm, right=3cm]{geometry}
\usepackage[latin1]{inputenc}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphics}
\usepackage{tabu}
\usepackage{breqn}
\usepackage[hang,flushmargin]{footmisc} 

\usepackage{amsmath,amsthm,amssymb,graphicx, multicol, array}
\usepackage{enumerate}
\usepackage{enumitem}
\usepackage{hyperref}
\usepackage{setspace}
\usepackage{xcolor}
\usepackage{currfile}

\usepackage{tabto}

\usepackage{pgfplots}
\pgfplotsset{width=10cm,compat=1.9}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}{Lemma}[section]
\newtheorem{conj}{Conjecture}[section]

\newcommand{\N}{\mathbb{N}}

\newcommand{\C}{\mathbb{C}}

\usepackage{fancyhdr}
\title{Elementary Proof Of The Siegel-Walfisz Theorem}
\date{}
\author{N. A. Carella}

\begin{document}
\thispagestyle{empty}
\date{}

\maketitle

\textbf{\textit{Abstract}:} This note offers an elementary proof of the \textit{Siegel-Walfisz theorem} for primes in arithmetic progressions. \let\thefootnote\relax\footnote{ \today \date{} \\
\textit{MSC2020}: Primary 11N13, Secondary 11N05. \\
\textit{Keywords}:  Prime in Arithmetic Progression, Prime Number Theorem, Siegel-Walfisz Theorem.}

\section{Introduction} \label{s4420}
The techniques used to prove the leading prime number theorems are classified as either complex analytic methods or elementary methods. The complex analytic methods rely on the principles of real and complex analysis, while the elementary methods do not rely on these techniques. Let $n\geq 1$ be an integer, and let the vonMangoldt function be defined by
\begin{equation} \label{eq773.52}
\Lambda(n)=
\begin{cases}
\log p & \text{if } n=p^m,\\ 
0 & \text{if } n\ne p^m.\\ 
\end{cases}
\end{equation}
The above notation $p^m\geq 2$, with $ m \in \N$, denotes a prime power. The prime number theorem claims that
\begin{equation} \label{eq4422.032}
\psi(x)=\sum_{p\leq x}\Lambda(n)\sim x,
\end{equation}
and 
\begin{equation} \label{eq4422.030}
\pi(x)=\sum_{p\leq x}1\sim \frac{x}{\log x}.
\end{equation} 
The first complex analytic proofs were given by delaVallee Poussin, and Hadamard, independently, (listed in alphabetical order). These proofs are based on the zerofree region $\mathcal{H}(\beta)=\{s\in \C: \mathcal{R}e(s) \geq 1 -\beta\}$ of the zeta function $\zeta(s)=\sum_{n \geq 1}n^{-s}$, where $s \in \C$ is a complex number, and $\beta=o(1)$ is a decreasing function of $s$, see \cite[Theorem 6.9]{MV07}. The first elementary proofs of the prime number theorem proofs were given by Erdos in \cite{EP49} and Selberg in \cite{SA49}, independently, (listed in alphabetical order). Several refinements and simplified versions of the elementary proofs appear in \cite{DS70}, \cite[Theorem 3.10]{EL85}, et alii. Although the elementary proofs are visibly independent of the zeta function, there is an indirect link to the zeta function and complex analysis, confer \cite[Section 2]{GR09} for an explication. \\

Let $a<q$ be a pair of small fixed integers such that $\gcd(a,q)=1$. The prime number theorem in arithmetic progression claims that
\begin{equation} \label{eq4422.092}
\psi(x,q,a)=\sum_{\substack{n\leq x\\ n\equiv a \bmod q}}\Lambda(n) \sim x.
\end{equation}
and 
\begin{equation} \label{eq4422.090}
\pi(x,q,a)=\sum_{\substack{p\leq x\\ p\equiv a \bmod q}}1 \sim \frac{x}{\varphi(q)\log x}.
\end{equation}
The first complex analytic proof was given by Dirichlet, see \cite[p.\ 49]{NW00}. The literature has many versions of the original complex analytic proof, see \cite[Theorem 2]{SP73}, \cite[Theorem 7.9]{EL85}, et cetera. The first elementary proof  was given by Selberg in \cite{SA49B}. An extended version appears in \cite{GR92}. Surveys of the early and new developments of the prime number theorem appear in \cite{BD96}, \cite{NW00}, \cite{FJ15}, and similar references. \\

An effective form of the prime number theorem in arithmetic progression, best known as the Siegel-Walfisz theorem, has the form
\begin{equation} \label{eq4422.096}
\psi(x,q,a)=\frac{x}{\varphi(q)} \left (1 +O\left (\frac{\varphi(q)}{(\log x)^B}\right )\right )  ,
\end{equation}
and 
\begin{equation} \label{eq4422.094}
\pi(x,q,a)=\frac{x}{\varphi(q)\log x} \left (1 +O\left (\frac{\varphi(q)}{(\log x)^{B-1}}\right )\right )  ,
\end{equation}
where $B>C+1$ is an arbitrary constant and $q=O\left((\log x)^C\right )$, respectively. Versions of the complex analytic proof appears in \cite[Theorem 8.8]{EL85}, \cite{HM72}, 
\cite[Corollary 11.19]{MV07}, and a discussion in \cite[p.\ 405]{FI10}. The standard proofs are based on the zerofree regions $\mathcal{H}(\beta_{\chi})=\{s\in \C: \mathcal{R}e(s) \geq 1 -\beta_{\chi}\}$ of the Dirichlet $L$-function $L(s,\chi)=\sum_{n \geq 1}\chi(n)n^{-s}$, where $s \in \C$ is a complex number, $\chi$ is a character modulo $q$, and $\beta_{\chi}=o(1)$ is a decreasing function of $s$. \\

The literature does not seem to have an elementary proof of the Siegel-Walfisz theorem. This note supplies an elementary proof of the Siegel-Walfisz theorem. 
	
	\begin{thm}	\label{thm4422.002} Let $x \geq 1$ be a large number, and let $a<q\ll(\log x)^{C}$ be a pair of relatively prime integers, and $C\geq 0$. Then,
\begin{enumerate} [font=\normalfont, label=(\roman*)]
\item  $\displaystyle \sum_{\substack{n \leq x\\ n \equiv a \bmod q}} \Lambda(n)=	\frac{x}{ \varphi(q)} \left (1+O\left(\frac{\varphi(q)}{(\log x)^{B}}\right )\right ),$
\item  $\displaystyle \sum_{\substack{n \leq x\\ n \equiv a \bmod q}}1=	\frac{x}{ \varphi(q)\log x} \left (1+O\left(\frac{\varphi(q)}{(\log x)^{B-1}}\right )\right ),$
\end{enumerate}	
where $B>C+1$ is an arbitrary constant, and $B>C+1$.
\end{thm}
The asymptotic formula is uniformly valid and nontrivial for all $C<B-1$. The restriction $q\ll(\log x)^{C}$ is imposed by the form of the error term in \eqref{eq4422.096}. \\

The simple proof in Section \ref{s4422} is a direct derivation from the Prime Number Theorem. This elementary method do not require any information on the zerofree region of the Dirichlet $L$-function $L(s,\chi)=\sum_{n \geq 1} \chi(n)n^{-s}$, and it is much simpler than the complex analytic methods.  \\

\section{Prime Numbers Theorems In Arithmetic Progressions}  \label{s4422}

\begin{proof} (i) Replace the identity for the vonMangoldt function, in Lemma \ref{lem22.002}, and rearrange it into main term and error term:  
\begin{eqnarray}\label{eq4422.012}
\sum_{\substack{n \leq x\\ n \equiv a \bmod q}} \Lambda(n)
&=&-\sum_{\substack{n \leq x\\ n \equiv a \bmod q}}\left ( \sum_{\substack{d \mid n\\ d <\sqrt{n}}}\mu(d)\log d +\sum_{\substack{d \mid n\\ d \leq\sqrt{n}}}\mu(n/d)\log (n/d) \right ) \\
&=&-\sum_{\substack{n \leq x\\ n \equiv a \bmod q,}}\sum_{\substack{d \mid n\\ d <\sqrt{n}}}\mu(d)\log d -\sum_{\substack{n \leq x\\ n \equiv a \bmod q,}}\sum_{\substack{d \mid n\\ d \leq\sqrt{n}}}\mu(n/d)\log (n/d)  \nonumber\\
	&=&-\sum_{d <\sqrt{x}}\mu(d)\log (d) \sum_{\substack{n \leq x\\ d \mid n \\ n \equiv a \bmod q}} 1-\sum_{d \leq\sqrt{x}} \sum_{\substack{n \geq x\\ d \mid n\\ n \equiv a \bmod q}} \mu(n/d)\log (n/d)\nonumber \\
&=&M(x)\quad +\quad E(x)\nonumber.
\end{eqnarray}
Substituting the evaluation in Lemma \ref{lem8822.100} for the main term, and the upper bound in Lemma \ref{lem8822.200} for the error term, return	
\begin{eqnarray}\label{eq4422.016}
\sum_{\substack{n \leq x\\ n \equiv a \bmod q}} \Lambda(n)
&=&M(x)\quad +\quad E(x) \\
&=&  \frac{x}{ \varphi(q)} \left (1+O\left(\frac{ \varphi(q)}{(\log x)^B}
	\right )\right )+ O\left(\frac{x}{ (\log x)^{B}}
	\right )\nonumber\\
	&=&\frac{x}{ \varphi(q)} \left (1+O\left( \frac{\varphi(q)}{(\log x)^{B}}
	\right )\right )\nonumber,
\end{eqnarray}
where $B>C+1$ is an arbitrary constant, and $q\ll(\log x)^{C}$. Clearly, this is nontrivial count for $B>C+1$.  (ii) Use partial summation.

\end{proof}
\section{Some Elementary Identities}  \label{s22.00}
These identities are sort of pre-hyperbola method technique. Nevertheless, these identities offer the same efficiency as the general hyperbola method, see \cite[Theorem 3.17]{AP98}, and \cite[Equation 2.9]{MV07}, et cetera.

\begin{lem} \label{lem22.001} If $n \geq 1$ is an integer, $\mu(n)$ is the Mobius function, and $\Lambda(n)$ is the vonMangoldt function, then,
\begin{equation}\label{22.100}
\Lambda(n)=-\sum_{d\mid n}\mu(d)\log d .
\end{equation}	
\end{lem}

\begin{proof} Let $\log n= \sum_{d \mid n} \log(d)= \sum_{d \mid n} \Lambda(d)$, and use the Mobius inversion formula to compute its inverse.
\end{proof}

\begin{lem} \label{lem22.002} If $n \geq 1$ is an integer, $\mu(n)$ is the Mobius function, and $\Lambda(n)$ is the vonMangoldt function, then,
\begin{equation}\label{22.102}
\Lambda(n)=-\sum_{\substack{d\mid n\\d < \sqrt{x}}}\mu(d)\log d -\sum_{\substack{d\mid n\\d \leq \sqrt{x}}}\mu(n/d)\log(n/d).
\end{equation}	
\end{lem}
\begin{proof} Employ the basic fact that the set of divisors $d \mid n$ of any integer $n\geq 2$ can be partition as a union of two disjoint subsets

\begin{equation}\label{22.120}
\{ d\mid n\}=\{ d\mid n: d < \sqrt{n}\} \cup \{ d\mid n: d \geq \sqrt{n}\}.
\end{equation}
Lastly, substitute $ \{ d\mid n: d \geq \sqrt{n}\}= \{ n/d\mid n: d \leq \sqrt{n}\}$.
\end{proof}
Other complicated versions of identity \eqref{22.102} are available in the literature. The best known among these is the Vaughan identity, see \cite[Section 2.6]{HG07}, and the literature for more information.
\section{Some Elementary Finite Sums}  \label{s8822}

\begin{lem} \label{lem8822.100} Let $a<q$ be a pair of small fixed integers such that $\gcd(a,q)=1$. If $x \geq 1$ is a large number, and $\mu(n)$ is the Mobius function, then,
\begin{enumerate} [font=\normalfont, label=(\roman*)]
\item  $\displaystyle -\sum_{d <\sqrt{x}}\mu(d)\log (d) \sum_{\substack{n \leq x\\ d \mid n \\ n \equiv a \bmod q}} 1= \frac{x}{ \varphi(q)} \left (1+O\left(\frac{ \varphi(q)}{(\log x)^B}
	\right )\right ),$ \\

where $B>1$ is a constant, unconditionally.
\item  $\displaystyle -\sum_{d <\sqrt{x}}\mu(d)\log (d) \sum_{\substack{n \leq x\\ d \mid n \\ n \equiv a \bmod q}} 1= \frac{x}{ \varphi(q)} \left (1+O\left(\frac{ \varphi(q)\log x}{ x^{1/2}}
	\right )\right ),$ \\

conditional on the RH.
\end{enumerate}	

\end{lem}

\begin{proof} (i) The local constraints
\begin{multicols}{2}
\begin{enumerate} [font=\normalfont, label=(\roman*)]
\item  $\displaystyle a\ne 0$,
\item  $\displaystyle \gcd(a,q)=1,$
\item  $\displaystyle n \equiv a \bmod q,$
\item$\displaystyle n \equiv 0 \bmod d,$
\end{enumerate}
\end{multicols}
imply that there is a unique $b\ne0$, $\gcd(b,q)=1$ such that $n \equiv b \bmod dq$. Moreover, since $n=dm$, it follows that $m\equiv c \bmod q$ for some $c\ne0$. Accordingly, the main term is equivalent to
\begin{eqnarray}\label{eq2099.003}
-\sum_{d <\sqrt{x}}\mu(d)\log (d) \sum_{\substack{n \leq x\\ d \mid n \\ n \equiv a \bmod q}} 1
&=&	-\sum_{d <\sqrt{x}}\mu(d)\log (d)\sum_{\substack{n \leq x\\ n \equiv b \bmod dq}} 1\\
&=&	-\sum_{d <\sqrt{x}}\mu(d)\log (d)\sum_{\substack{m \leq x/d\\ m \equiv c \bmod q}} 1\nonumber.
\end{eqnarray}
Let $\{z\}\in (0,1)$ denotes the fractional part function. Then, the main term has the asymptotic expression
\begin{eqnarray}\label{eq2099.004}
-\sum_{d <\sqrt{x}}\mu(d)\log (d)\sum_{\substack{m \leq x/d\\ m \equiv c \bmod q}} 1&=&-\frac{1}{ \varphi(q)}\sum_{d <\sqrt{x}}\mu(d)\log (d)\left ( \frac{x}{d }-\left \{ \frac{x}{d }\right \}\right )\\
&=&-\frac{x}{ \varphi(q)} \sum_{d <\sqrt{x}}\frac{\mu(d)\log (d)}{d} -\sum_{d <\sqrt{x}}\mu(d)\log (d)\left \{ \frac{x}{d }\right \}\nonumber \\
&=&-\frac{x}{ \varphi(q)} \sum_{d <\sqrt{x}}\frac{\mu(d)\log (d)}{d} +O\left( \sqrt{x} (\log x)\right )\nonumber .
\end{eqnarray}
The partial sum in the third line of \eqref{eq2099.004} converges to a constant:
	\begin{equation}\label{eq2099.006}
	\sum_{n <\sqrt{x}}\frac{\mu(n)\log (n)}{n}=\sum_{n \geq 1}\frac{\mu(n)\log (n)}{n}-\sum_{d \geq \sqrt{x}}\frac{\mu(n)\log (n)}{n}=-1+
	O\left(\frac{ 1}{(\log x)^B}
	\right ),
	\end{equation}
where $B>1$ is a constant, see Lemma \ref{lem8822.300} for the detailed calculation. Merging \eqref{eq2099.004} and \eqref{eq2099.006} yield
\begin{eqnarray}\label{eq2099.008}
-\sum_{d <\sqrt{x}}\mu(d)\log (d)\sum_{\substack{m \leq x/d\\ m \equiv c \bmod q}} 1
&=&-\frac{x}{ \varphi(q)} \sum_{d <\sqrt{x}}\frac{\mu(d)\log (d)}{d} +O\left( \sqrt{x} (\log x)\right ) \\
&=&\frac{x}{ \varphi(q)} \left (1+O\left(\frac{  \varphi(q)}{(\log x)^B}
	\right )\right )\nonumber.
\end{eqnarray}
(ii) The partial sum in the third line of \eqref{eq2099.004} converges to a constant:
	\begin{equation}\label{eq2099.010}
	\sum_{n <\sqrt{x}}\frac{\mu(n)\log (n)}{n}=\sum_{n \geq 1}\frac{\mu(n)\log (n)}{n}-\sum_{d \geq \sqrt{x}}\frac{\mu(n)\log (n)}{n}=-1+
	O\left(\frac{ (\log x)^2}{ x^{1/2}}
	\right ),
	\end{equation}
Merging \eqref{eq2099.004} and \eqref{eq2099.008} to complete the proof.
\end{proof}

\begin{lem} \label{lem8822.200} Let $a<q$ be a pair of small fixed integers such that $\gcd(a,q)=1$. If $x \geq 1$ is a large number, and $\mu(n)$ is the Mobius function, then,
\begin{enumerate}[font=\normalfont, label=(\roman*)]
\item $\displaystyle-\sum_{d \leq\sqrt{x}} \sum_{\substack{n \leq x\\ d \mid n\\ n \equiv a \bmod q}} \mu(n/d)\log (n/d)= O\left(\frac{x}{ (\log x)^{B}}
	\right ),$\\

where $B>1$ is a constant, unconditionally.
    
 \item $\displaystyle -\sum_{d \leq\sqrt{x}} \sum_{\substack{n \leq x\\ d \mid n\\ n \equiv a \bmod q}} \mu(n/d)\log (n/d)= O\left(x^{1/2+\varepsilon}\right ),$\\

conditional on the RH for any small number $\varepsilon>0$.
 \end{enumerate}
\end{lem}

\begin{proof}(i) Let $n=dm\leq x$. The local constraints
\begin{multicols}{2}
\begin{enumerate} [font=\normalfont, label=(\roman*)]
\item  $\displaystyle a\ne 0$,
\item  $\displaystyle \gcd(a,q)=1,$
\item  $\displaystyle n \equiv a \bmod q,$
\item$\displaystyle n \equiv 0 \bmod d,$
\end{enumerate}
\end{multicols}
imply that there is a unique $b\ne0$, $\gcd(b,q)=1$ such that $n \equiv b \bmod dq$. Moreover, (iii) and (iv) imply that $m\equiv c \bmod q$ for some $c\ne0$. Thus, the error term has the asymptotic
	\begin{eqnarray}\label{eq2099.008}
	-\sum_{d \leq\sqrt{x}} \sum_{\substack{n \leq x\\ d \mid n\\ n \equiv a \bmod q}} \mu(n/d)\log (n/d)
	&=&-\sum_{d \leq\sqrt{x}} \sum_{\substack{n \leq x\\ n \equiv b \bmod dq}} \mu(n/d)\log (n/d)\\
	&=&-\sum_{d \leq\sqrt{x}} \sum_{\substack{m \leq x/d\\ m \equiv c \bmod q}} \mu(m)\log (m)\nonumber. 
	\end{eqnarray}
Applying Theorem \ref{thm8822.459} to the last finite sum in \eqref{eq2099.008} yield the upper bound
\begin{eqnarray}\label{eq2099.010}
-\sum_{d \leq\sqrt{x}} \sum_{\substack{m \leq x/d\\ m \equiv c \bmod q}} \mu(m)\log (m)
&=&O\left(\frac{x}{(\log x)^{D}}
\sum_{d \leq\sqrt{x}} \frac{1}{d}\right )\\
&=&O\left(\frac{x}{ (\log x)^{D-1}}	\right )\nonumber, 
\end{eqnarray}
where $a<q$ be a pair of small fixed integers such that $\gcd(a,q)=1$, and $B>D+1$ is an arbitrary constant.	
\end{proof}

\section{Some Elementary Foundation}  \label{s8822}
The relations
\begin{equation}\label{eq8822.300}
\sum_{n \geq1} \frac{ \mu(n)}{n}=0\qquad \text{ and } \qquad \sum_{n \geq1} \frac{ \mu(n)\log (n)}{n}=-1
\end{equation}
are known to be equivalent to the prime number theorem, as claimed in \cite[p.\ 78]{EL85}, \cite[p.\ 285]{NW00}, et alii. An elementary proof of the second one is given \textit{in situ}.
\begin{lem} \label{lem8822.300} Let $\mu(n)$ be the Mobius function, then,
\begin{equation}\label{eq8822.302}
\frac{\zeta^{\prime}(1)}{\zeta(1)^2}=\sum_{n \geq1} \frac{ \mu(n)\log (n)}{n}=-1.
\end{equation}	
\end{lem}

\begin{proof} Let $1/\zeta(s)=\sum_{n \geq 1}\mu(n)n^{-s}$. The derivative at $s=1$ can be computed using the Taylor series 
\begin{equation}\label{eq8822.303}
\zeta(s)= \frac{ 1}{s-1}+c_0+c_1(s-1)+c_2(s-1)^2+\cdots,
\end{equation}
of the zeta function at $s=1$, see \cite[Equation 25.2.4]{DLMF}, and a derivation in \cite{BF88}. Specifically,
\begin{equation}\label{eq8822.305}
\frac{\zeta^{\prime}(s)}{\zeta(s)^2}= \frac{\frac{- 1}{(s-1)^2}+c_1+c_2(s-1)+\cdots}{\left (\frac{ 1}{s-1}+c_0+c_1(s-1)+c_2(s-1)^2+\cdots\right )^2}.
\end{equation}
Taking the limit at $s=1$ yields
\begin{equation}\label{eq8822.307}
\lim_{s\to 1} \frac{(s-1)^2}{(s-1)^2}\frac{\zeta^{\prime}(s)}{\zeta(s)^2}= -1.
\end{equation}
\end{proof}
There are other means of evaluating the derivative, see \cite[Exersice 16, p.\ 184]{MV07}. \\

The basic estimates for finite sums of the forms $\sum_{n \leq x} \mu(n)\log (n)$ used in Lemma \ref{lem8822.200} are derive from the standard results given below.
\begin{thm} \label{thm8822.459} Let $a<q$ be a pair of small fixed integers such that $\gcd(a,q)=1$, and  let $\mu(n)$ be the Mobius function. Then, 
\begin{enumerate}[font=\normalfont, label=(\roman*)]
\item $\displaystyle \sum_{\substack{n \leq x\\ n \equiv a \bmod q}} \mu(n)=O\left(\frac{x}{\log^D x} \right ),$ \tabto{6cm} unconditionally, for any constant $D>0$,
    
 \item $\displaystyle \sum_{\substack{n \leq x\\ n \equiv a \bmod q}} \mu(n)=O\left(x^{1/2+\varepsilon} \right ),$ \tabto{6cm} conditional on the RH, for any constant $\varepsilon>0$.
 \end{enumerate}
\end{thm}

\begin{proof}Confer \cite[p.\ 182]{MV07}, and similar references.
\end{proof}
\section{ Conditional Results And Conjecture}  \label{s2030}
Under the generalized Riemann hypothesis, the prime number theorem for the number of primes $p\leq x$ in the arithmetic progression 

\begin{equation} \label{eq2030.05}
\mathcal{A}(q,a)=\{p=qn+a: n \geq 1\},
\end{equation}
specifies the asymptotic formulas
\begin{equation} \label{eq2030.00}
\psi(x,q,a)=\frac{x}{\varphi(q)}  +O\left (x^{1/2}(\log x)^2\right ),
\end{equation}
and
\begin{equation} \label{eq2030.00}
\pi(x,q,a)=\frac{x}{\varphi(q)\log x}  +O\left (x^{1/2}\log x\right ).
\end{equation}
This is uniformly nontrivial for $q\ll x^{1/2}/\log(x)^C$. This restriction is imposed by the form of the error term in \eqref{eq2030.00}, more information appear in \cite[p.\ 419]{IK04}. \\

Under the Riemann hypothesis, both of the expressions \eqref{eq2030.05} and \eqref{eq2030.05} can be derived using the same elementary techniques demonstrated in Sections \ref{s4422} to \ref{s8822}, and it is independent of the zerofree regions of the corresponding $L$-function.  The precise statements are as follow.

\begin{thm} \label{thm2030.200} Assume the RH. Let $1\leq a<q\ll x^{1/2}/\log(x)^C$ be a pair of small fixed integers such that $\gcd(a,q)=1$, and $C\geq 2$ is a constant. Then, 
\begin{enumerate}[font=\normalfont, label=(\roman*)]
\item $\displaystyle \psi(x,q,a)=\frac{x}{\varphi(q)}  +O\left (x^{1/2}(\log x)^2\right ),$
    
 \item $\displaystyle \pi(x,q,a)=\frac{x}{\varphi(q)\log x}  +O\left (x^{1/2}\log x\right ).$
 \end{enumerate}
\end{thm}

\begin{proof} (i) Proceed as in \eqref{eq4422.012}, and \eqref{eq4422.016}. Substituting the conditional evaluation in Lemma \ref{lem8822.100} for the main term, and the conditional upper bound in Lemma \ref{lem8822.200} for the error term, return	
\begin{eqnarray}\label{eq2030.016}
\sum_{\substack{n \leq x\\ n \equiv a \bmod q}} \Lambda(n)
&=&M(x)\quad +\quad E(x) \\
&=&  \frac{x}{ \varphi(q)} \left (1+O\left(\frac{ \varphi(q)\log x}{x^{1/2}}
	\right )\right )+ O\left(x^{1/2} (\log x)^{2}
	\right )\nonumber\\
	&=&\frac{x}{ \varphi(q)} \left (1+O\left(\frac{ \varphi(q)(\log x)^{2}}{x^{1/2}}
	\right )\right )\nonumber.
\end{eqnarray}
(ii) Use partial summation.
\end{proof}

The Montgomery conjecture, stated below, does not seem to have similar elementary proof derived directly from the RH.
\begin{conj}\label{conj2099.35} {\normalfont (\cite[Conjecture 13.9]{MV07}} Let $a<q$ be integers, $\gcd(a,q)=1$, and $q \leq x$. Then,
\begin{enumerate} [font=\normalfont, label=(\roman*)]
\item  $\displaystyle \psi(x,q,a)=\frac{x}{\varphi(q)}+O\left ( \frac{x^{1/2+\varepsilon}}{q^{1/2}}\right ).$
\item  $\displaystyle \pi(x,q,a)=\frac{x}{\varphi(q)\log x}+O\left ( \frac{x^{1/2+\varepsilon}}{q^{1/2}}\right ).$
\end{enumerate}	

\end{conj}
The Montgomery conjecture for primes in arithmetic progression, recorded above, extends the range of moduli to nearly all $q<x$, but not all $q\leq x$. The advanced theory for equidistributions and oscillations in the prime number theorems was pioneered and developed in a series of papers, \cite{MH88}, \cite{FG91}, et cetera.


\currfilename.\\
\end{document}